MalayaJanae:

Evaluate the expression 3.14(a2 + ab) when a = 4 and b = 3. (Input decimals only, such as 12.71, as the answer.)

ZombieLover:

3.14(2(4)+(4)(2))=3.14(8+8)=3.14(16)=50.24

MalayaJanae:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @ZombieLover 3.14(2(4)+(4)(2))=3.14(8+8)=3.14(16)=50.24 \(\color{#0cbb34}{\text{End of Quote}}\) Nice Job! Keep up the good work!